OpenStudy (anonymous):

solve for x..... e^2x=1

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

because anything to the power of 0 is 1

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

\[b^0=1\] by definition

OpenStudy (anonymous):

it isn't to the power of zero

OpenStudy (anonymous):

you have \[e^{2x}=1\] yes?

OpenStudy (anonymous):

the equation is e to the power of 2x which equals 1

OpenStudy (anonymous):

so i don't see how you get that x would be zero

OpenStudy (anonymous):

right. so what power do you raise any number to to give 1?

OpenStudy (anonymous):

ohhh zero?

OpenStudy (anonymous):

yes so you know \[2x=0\]

OpenStudy (anonymous):

and therefore x = ...

OpenStudy (anonymous):

ohhh okay thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

do you think i could askyou another question?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

ln(x^2)=ln(x+2)

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\ln(x)\] is a "one to one" function which is a fancy way of saying that if \[\ln(a)=\ln(b)\] then \[a=b\]

OpenStudy (anonymous):

ohhh so x^2 equals x+2?

OpenStudy (anonymous):

you are given that \[\ln(x^2)=\ln(x+2)\] and the only way that is true is if \[x^2=x+2\]

OpenStudy (anonymous):

just be careful when you solve.

OpenStudy (anonymous):

\[x^2-x-2=0\] \[(x-2)(x+1)=0\] \[x=2\] or \[x=-1\]

OpenStudy (anonymous):

yes you are right. solve that

OpenStudy (anonymous):

now we check the answers, because you cannot take the log of a negative number

OpenStudy (anonymous):

but that is ok here because if x = -1 you get \[\ln((-1)^2)=\ln(1)=0\] and also \[\ln(-1+2)=\ln(1)=0\] so -1 works also 2 works because \[\ln(2^2)=\ln(2+2)=\ln(4)\]

OpenStudy (anonymous):

ok so can i keep asking you questions or is it annoying?

OpenStudy (anonymous):

one more then i have to go

OpenStudy (anonymous):

cos^2x+2cosx=0

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

factor to get \[\cos(x)(\cos(x)+2)=0\]

OpenStudy (anonymous):

first one give \[\cos(x)=0\] so \[x=\frac{\pi}{2}\] or \[x=\frac{3\pi}{2}\]

OpenStudy (anonymous):

second one give \[\cos(x)=-2\] which is not possible since cosine is between -1 and 1

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

dont go!! haha

OpenStudy (anonymous):

one more?